OpenStudy (anonymous):

SHOW WORK! 8x+5=2x-7 and this one 4y-2+3y=15+y-5

OpenStudy (callisto):

8x+5=2x-7 8x-2x = -7-5 6x = -12 x=-2

OpenStudy (callisto):

4y-2+3y=15+y-5 7y -2 = 10+y 7y - y = 10+2 6y = 12 y=2